# SegAnyPET: Universal Promptable Segmentation from Positron Emission Tomography Images

### v3

@@ Abstract
Positron Emission Tomography (PET) is a powerful molecular imaging tool that plays a crucial role in modern medical diagnostics by visualizing radio-tracer distribution to reveal physiological processes. Accurate organ segmentation from PET images is essential for comprehensive multi-systemic analysis of interactions between different organs and pathologies. Existing segmentation methods are limited by insufficient annotation data and varying levels of annotation, resulting in weak generalization ability and difficulty in clinical application. Recent developments in segmentation foundation models have shown superior versatility across diverse segmentation tasks. Despite the efforts of medical adaptations, these works primarily focus on structural medical images with detailed physiological structural information and exhibit limited generalization performance on molecular PET imaging. In this paper, we collect and construct PETS-5k, the largest PET segmentation dataset to date, comprising 5,731 three-dimensional whole-body PET images and encompassing over 1.3M 2D images. Based on the established dataset, we develop SegAnyPET, a modality-specific 3D foundation model for universal promptable segmentation from PET images. To issue the challenge of discrepant annotation quality, we adopt a cross prompting confident learning (CPCL) strategy with an uncertainty-guided self-rectification process to robustly learn segmentation from high-quality labeled data and low-quality noisy labeled data for promptable segmentation. Experimental results demonstrate that SegAnyPET can segment seen and unseen target organs using only one or a few prompt points, outperforming state-of-the-art foundation models and task-specific fully supervised models with higher accuracy and strong generalization ability for universal segmentation. As the first foundation model for PET images, we believe that SegAnyPET will advance the applications to various downstream tasks for molecular imaging.  [^1] The model and code will be publicly available at the project page.
@@ 1 Introduction
Medical imaging technology has become an indispensable tool in modern clinical diagnosis [14, 19]. The steady development of increasingly sophisticated imaging techniques has led to a dramatic increase in patient treatment outcomes, especially when diseases can only be observed internally within the patient’s body [18]. In addition to commonly used medical modalities such as computed tomography (CT) and magnetic resonance imaging (MRI), which can provide detailed physiological structural information, positron emission tomography (PET) has also been widely applied in medical examinations as an emerging molecular imaging modality [31]. The purpose of PET imaging is to reveal the ongoing metabolic processes within the body of a patient. The radioactive decay of injected tracer emits positrons, leading to gamma photons, which are produced by annihilation and captured by detectors surrounding the patient to create a 3D map of metabolic activity. As the most widely used tracer, fluorodeoxyglucose (FDG) is used to assess local glucose uptake and evaluate organ metabolism [29] and presence of tumor metastasis [4], so as to monitor progress during the treatment process. Accurate organ segmentation from PET images is essential for comprehensive multi-systemic analysis of interactions between different organs and pathologies [35, 33, 45]. However, due to the low resolution nature and the partial volume effects, PET images exhibit lower image quality and have less distinct boundaries compared to other structural images like CT (see Fig. 1), which presents a significant challenge for accurate segmentation.
@@
With the unprecedented developments in deep learning, deep neural networks have been widely applied and achieved great success in many medical image segmentation tasks [30, 9]. However, the availability of PET images and annotations is far behind that of CT and MRI [20, 11]. A significant obstacle lies in the high costs associated with acquisition and annotation of PET images, and the varying levels of annotation due to the low image quality and the presence of partial volume effects. Although there are a few publicly available PET datasets [4, 27], they only focus on limited targets and do not cover these clinically important anatomic structures. In addition, existing deep models are often tailored for specific tasks, which require task-specific training data for model development and lack the capacity for generalization across different tasks. When adapted to new segmentation task, additional training with annotated data is required.
@@
The emergence of pre-trained foundation models has sparked a new era due to their remarkable generalization abilities across a wide range of downstream tasks [24, 37]. For image segmentation, the introduction of the Segment Anything Model (SAM) [15] has gained massive attention as a promptable foundation model capable of generating fine-grade segmentation masks to unseen targets using positional prompts like points or bounding boxes. However, the original SAM is trained on natural images characterized by strong edge information, which differs significantly from medical images. When directly applying SAM without any adaptation, its performance significantly left behind task-specific segmentation models [43, 23]. Despite the attempts of fine-tuning and adapting SAM for medical images [21, 5], these works mainly focus on modalities like CT and MRI with detailed physiological structural information, while struggling to generalize to molecular PET imaging.
@@
To bridge this gap, we collect and construct a large-scale PET dataset named PETS-5k with voxel-level multi-organ annotation. The dataset consists of 5,731 three-dimensional whole-body PET images, encompassing over 1.3M 2D images, which is the largest PET dataset to date. Based on the empirical analysis, we found that existing foundation models show poor generalization performance on PET images. To enable universal promptable segmentation from PET images, we present SegAnyPET, a modality-specific foundation model that can be efficiently and robustly adapted for universal promptable segmentation from PET images. Other than the original 2D architecture of SAM, SegAnyPET utilizes a 3D architecture to fully utilize the inter-slice context information of 3D PET volumes. To issue the challenge of discrepant annotation quality, we adopt a cross prompting confident learning (CPCL) strategy with an uncertainty-guided self-rectification process to efficiently learn from dataset with varying annotation quality for model training. The strategy does not require network modification and can be easily adapted to other promptable foundation models. Extensive experiments demonstrate that SegAnyPET can correctly segment both seen and unseen targets using only one or few prompt points and outperform state-of-the-art foundation models and task-specific fully supervised models by a large margin, underpinning its general-purpose segmentation ability. Our contribution can be summarized as follows:
@@
  • A cross prompting confident learning (CPCL) strategy for noisy-robust learning to efficiently learn from high-quality labeled data and low-quality noisy labeled data.
@@ 2 Related Work
Vision Foundation Models. Foundation models represent a rapidly expanding field in artificial intelligence research, focusing on developing large-scale, general-purpose models with capabilities applicable across various domains and applications. Pioneering vision foundation models have been primarily based on pre-training methods such as CLIP [28] and ALIGN [12] leverage contrastive learning techniques to train both text and image encoders. However, these models primarily excel in tasks that involve mapping images to text, such as classification. For foundation model of image segmentation, the Segment Anything Model (SAM) [15] represents a new paradigm of image segmentation with universal segmentation ability across different tasks and demonstrates strong potential in addressing a wide range of downstream tasks such as medical image analysis [40, 24].
Foundation Models for Medical Image Segmentation. As a crucial branch of image segmentation, recent studies have explored the generalization ability of SAM to biomedical image segmentation tasks of different modalities and targets [23, 7, 22], while some research also focuses on fine-tuning SAM [21] or adapting SAM-like foundation models [34, 3, 38] on large-scale medical datasets. These studies indicate that SAM has limited generalization ability when directly applied to medical image segmentation due to the significant distribution gap, and fine-tuning SAM or training promptable segmentation foundation models on medical datasets can improve the unsatisfactory performance [43]. Other than developing promptable foundation models for medical imaging, another line of research aims to extend the promptable interactive segmentation manner into end-to-end fully automatic segmentation by auto-prompting [16, 41] or synergy in training supervised segmentation models [44, 17]. However, due to the scarcity of public datasets, existing adaptions mostly focus on structural medical images, while ignoring molecular PET images.
Annotation-Efficient Learning. Despite the outstanding performance in many medical image segmentation tasks [9, 36], the success of existing deep neural networks heavily relies on the massive radiologist-examined labeled data. Obtaining pixel-wise annotation is time-consuming and expertise-demanding, which motivates annotation-efficient learning strategies to train deep models with limited annotations [2, 13, 32]. Despite recent efforts of large-scale medical datasets [39, 1], these datasets are the integration of existing public datasets with different annotation quality. While some labels may be noisy with mistakes and could potentially hinder the learning procedure, it is important to develop noise-robust learning strategies to issue this challenge.
@@ 3 Methodology
6:        Input prompt $p_{i}$Leftarrow PromptGenerate($Y$, $\hat{Y}_{i-1}$)
@@
3.1 Dataset Construction
To develop a foundation model for universal segmentation from PET images, we collect and construct a large-scale PET dataset with 5,731 three-dimensional whole-body PET images, which is the largest PET dataset to date. In contrast to existing PET datasets with limited annotated targets, our work focuses on addressing a broader spectrum of multi-organ segmentation from whole-body images to facilitate representation learning for universal segmentation. All images are preliminarily annotated by AI and junior annotators to form the LQ set. A subset is then checked and refined by senior experts with quality control, serving as the HQ set. More information of the dataset are shown in the Appendix.
@@ 3.2 Network Architecture
Motivated by the recent advance of segmentation foundation models for natural images [15], we reformulate a holistic 3D structure in SegAnyPET to capture the spatial information directly from volumetric images for universal segmentation. Specifically, SegAnyPET consists of three main components, including an image encoder $\mathcal{F}_{IE}$, a prompt encoder $\mathcal{F}_{PE}$, and a mask decoder $\mathcal{F}_{D}$. The image encoder aims to transform input images into discrete embeddings, and the prompt encoder converts input prompts into compact embeddings by combining fixed positional encoding and adaptable prompt-specific embeddings. After that, the mask decoder receives the extracted information from both the image encoder and the prompt encoder, which is then up-sampled and passes through a multi-layer perception to generate the output segmentation masks. The processed feature map is then up-sampled and passes through a multi-layer perception to generate the output segmentation masks. Details of the network components are shown in the Appendix. Given an input image $X$ and spatial prompt $p$, the output segmentation mask $\hat{Y}$ can be formally expressed as follows:
$\hat{Y}_{i}=\left\{\begin{array}[]{lllll}\mathbf{O}_{H\times W\times D},&i=0\\ \\ \mathcal{F}_{D}(\mathcal{F}_{IE}(X),\mathcal{F}_{PE}(p_{i},\hat{Y}_{i-1})),&i\geq 1\\ \end{array}\right.$ (1)
where $\mathbf{O}_{H\times W\times D}$ is an all-zero matrix as the pre-defined initial segmentation mask. $p_{i}$ represents the $i_{th}$ prompt point in the training loop, where new point prompts are generated from error regions based on the previous segmentation mask to simulate the manual interactive segmentation workflow.
@@
SegResNet (HQ) [25] Auto 92.28 85.56 82.94 88.21 86.55 87.11
SwinUNETR (HQ) [6] Auto 93.03 89.23 83.62 89.03 87.93 88.57
nnUNet (HQ) [9] Auto 93.16 90.12 86.60 90.96 88.32 89.83
STUNet (HQ) [8] Auto 93.15 90.11 85.60 90.38 88.83 89.61
SegResNet (All) [25] Auto 91.43 86.29 85.73 88.88 86.45 87.76
SwinUNETR (All) [6] Auto 91.15 84.98 84.32 88.13 84.89 86.69
nnUNet (All) [9] Auto 93.74 88.21 88.26 90.50 88.30 89.80
STUNet (All) [8] Auto 93.45 87.97 87.67 90.21 88.06 89.47
Table 1: Comparison of DSC performance with state-of-the-art general-propose foundation models for interactive segmentation and trainable task-specific models for automatic segmentation from PET images. N denotes the count of slices containing the target object (N ranges from 20 to 50 in our task).
@@ 3.3 Cross Prompting Confident Learning
Due to the dependency of domain knowledge for annotating medical images, acquiring large amounts of high-quality labeled data is infeasible. A prevalent approach to mitigating this issue is collecting additional labeled data with varying annotation qualities like crowdsourcing from non-experts or using model-generated labels. Given that some annotations may be fraught with noise or inaccuracies, which could impede the learning process, it is imperative to devise noise-robust learning strategies to efficiently learn from both HQ Set containing expert-examined high-quality annotations and LQ Set containing low-quality annotations with possible mistakes. This challenge is particularly pronounced for PET images due to the presence of partial volume artifacts and low image quality, and directly learning from these low-quality annotations may provide undesirable guidance. To issue the challenge of discrepant annotation quality, we adopt a cross prompting confident learning (CPCL) strategy. Figure 3 depicts the workflow of CPCL. The utilization of LQ set can be mainly divided into the following two parts.
@@ 3.4 Overall Training Procedure
Aorta Lung-LL Lung-LR Lung-UL Lung-UR Lung-MR Prostate Avg
@@
Table 2: Generalization performance to training invisible organs with comparison to state-of-the-art state-of-the-art general-propose foundation models for interactive segmentation from PET images.
4 Experiment & Results
4.1 Experimental Setup
@@
The proposed PETS-5k dataset consists of 5,731 three-dimensional whole-body PET images is used for training and internal validation. Among the dataset, a subset of 100 images is with high-quality annotations verified by domain experts, which is split into 40 cases serving as HQ training set and 60 cases serving as internal test set. The remaining images with possible mislabeled or unlabeled pixels serving as the LQ training set. To evaluate the generalization ability for universal segmentation, we annotate 7 additional training invisible organs from the internal test set. We further annotate a small subset of 100 cases with all 12 testing organs from the AutoPET dataset [4] for external validation, entitled AutoPET-Organ. Following the setting for promptable segmentation, the input prompts are simulated based on the ground truth mask with random perturbations. We compare proposed SegAnyPET with state-of-the-art segmentation foundation models [15, 21, 34] and training-based task-specific models [25, 9, 6, 8]. More information of the implementation details are shown in the Appendix.
4.2 Experimental Results

### v4

@@ Abstract
Positron Emission Tomography (PET) imaging plays a crucial role in modern medical diagnostics by revealing the metabolic processes within a patient’s body, which is essential for quantification of therapy response and monitoring treatment progress. However, the segmentation of PET images presents unique challenges due to their lower contrast and less distinct boundaries compared to other structural medical modalities. Recent developments in segmentation foundation models have shown superior versatility across diverse natural image segmentation tasks. Despite the efforts of medical adaptations, these works primarily focus on structural medical images with detailed physiological structural information and exhibit poor generalization ability when adapted to molecular PET imaging. In this paper, we collect and construct PETS-5k, the largest PET segmentation dataset to date, comprising 5,731 three-dimensional whole-body PET images and encompassing over 1.3M 2D images. Based on the established dataset, we develop SegAnyPET, a modality-specific 3D foundation model for universal promptable segmentation from PET images. To issue the challenge of discrepant annotation quality of PET images, we adopt a cross prompting confident learning (CPCL) strategy with an uncertainty-guided self-rectification process to robustly learn segmentation from high-quality labeled data and low-quality noisy labeled data. Experimental results demonstrate that SegAnyPET can correctly segment seen and unseen targets using only one or a few prompt points, outperforming state-of-the-art foundation models and task-specific fully supervised models with higher accuracy and strong generalization ability for universal segmentation. As the first foundation model for PET images, we believe that SegAnyPET will advance the applications to various downstream tasks for molecular imaging. [^1] The model and code will be publicly available at the project page.
* equal contribution. †corresponding authors.
@@ 1 Introduction
Medical imaging technology has become an indispensable tool in modern clinical diagnosis [14, 19]. The steady development of increasingly sophisticated imaging techniques has led to a dramatic increase in patient treatment outcomes, especially when diseases can only be observed internally within the patient’s body [18]. Segmentation of targets such as organs and tumors from medical images is one of the most representative and comprehensive research topics in both the computer vision and medical image analysis communities [hesamian2019deep]. Accurate segmentation can provide reliable volumetric and shape information on target structures and assist in many clinical applications [lalande2021deep, 20, zhang2024nasalseg].
In addition to commonly used medical modalities such as computed tomography (CT) and magnetic resonance imaging (MRI), which can provide detailed physiological structural information, positron emission tomography (PET) has also been widely applied in medical examinations as an emerging molecular imaging modality [31]. The purpose of PET imaging is to reveal the ongoing metabolic processes within the body of a patient. A radioactive tracer is injected into the patient, typically containing the radioactive isotope fluorine-18 synthesized fluorodeoxyglucose (18F-FDG). The radioactive decay emits positrons, leading to gamma photons, which are produced by annihilation and captured by detectors surrounding the patient to create a 3D map of metabolic activity. In FDG-PET examinations, the tracer is used to assess local glucose uptake and evaluate organ metabolism [29, 35] and presence of tumor metastasis [4], so as to monitor progress during the treatment process. However, due to the low resolution nature and the partial volume effects, PET images exhibit lower image quality and have less distinct boundaries compared to other structural images such as CT scans (see Fig. 1), which presents a significant challenge for accurate segmentation.
@@
With the unprecedented developments in deep learning, deep neural networks have been widely applied and achieved great success in medical image segmentation [30, 9]. However, existing deep models are often tailored for specific tasks, which require task-specific training data for model development and lack the capacity for generalization across different tasks. The emergence of pre-trained foundation models has sparked a new era due to their remarkable generalization abilities across a wide range of downstream tasks [24, 37]. For image segmentation, the introduction of the Segment Anything Model (SAM) [15] has gained massive attention as a promptable foundation model capable of generating fine-grade segmentation masks to unseen targets using positional prompts like points or bounding boxes. However, the original SAM is trained on natural images characterized by strong edge information, which differs significantly from medical images, especially PET images that exhibit low contrast and weak boundaries. When directly applying SAM without any adaptation, its performance significantly left behind task-specific segmentation models [43, 23]. Despite the attempts of fine-tuning and adapting SAM for medical images [21, 5], these works mostly focus on CT and MRI images with detailed physiological structural information, ignoring the PET imaging. A significant obstacle lies in the scarcity of large-scale annotated PET datasets, which is attributable to the high costs associated with acquisition and annotation. Existing PET datasets [4, 27] are relatively small and only focus on the segmentation of limited targets. Besides, the low image quality and the presence of partial volume effects complicate the annotation of segmentation targets from PET images, potentially leading to inconsistent annotation quality.
@@
To overcome this challenge, we collect and construct a large-scale PET segmentation dataset named PETS-5k. The dataset consists of 5,731 three-dimensional whole-body PET images, encompassing over 1.3M 2D images, which is the largest PET dataset to date. Based on the empirical analysis, we found that existing foundation models show poor generalization performance on PET images. To enable universal promptable segmentation from PET images, we propose SegAnyPET, a modality-specific foundation model that can be efficiently and robustly adapted to segment anything from PET images. Other than the original 2D architecture of SAM, we reformulate a 3D architecture to fully utilize the inter-slice context information of 3D PET images. To issue the challenge of discrepant annotation quality, we adopt a cross prompting confident learning (CPCL) strategy with an uncertainty-guided self-rectification process to efficiently learn from dataset with varying annotation quality for model training. The strategy does not require network modification and can be easily adapted to other promptable foundation models. Extensive experiments demonstrate that SegAnyPET can correctly segment both seen and unseen targets using only one or few prompt points and outperform state-of-the-art foundation models and task-specific fully supervised models by a large margin, underpinning its general-purpose segmentation ability. Our contribution can be summarized as follows:
@@
  • A cross prompting confident learning (CPCL) strategy for noisy-robust learning to efficiently learn from high-quality labeled data and low-quality noisy labeled data for model training.
@@ 2 Related Work
Vision Foundation Models. Foundation models represent a rapidly expanding field in artificial intelligence research, focusing on developing large-scale, general-purpose models with capabilities applicable across various domains and applications. Pioneering vision foundation models have been primarily based on pre-training methods such as CLIP [28] and ALIGN [12] leverage contrastive learning techniques to train both text and image encoders. However, these models primarily excel in tasks that involve mapping images to text, such as classification. For foundation model of image segmentation, the Segment Anything Model (SAM) [15] represents a new paradigm of image segmentation with universal segmentation ability across different tasks and demonstrates strong potential in addressing a wide range of downstream tasks like medical image analysis [40, 24].
Foundation Models for Medical Image Segmentation. As a crucial branch of image segmentation, recent studies have explored the application of SAM to the segmentation tasks from multi-modal multi-target biomedical images [23, 7, 22], while some research also focuses on fine-tuning SAM [21] or adapting SAM-like foundation models [34, du2023segvol, 38] on large-scale medical datasets. These studies indicate that SAM has limited generalization ability when directly applied to medical image segmentation due to the significant distribution gap, and fine-tuning SAM on medical datasets can improve the unsatisfactory performance [43]. In addition to the adaptation of model architectures, another line of research aims to extend the original promptable interactive segmentation method into end-to-end fully automatic segmentation by auto-prompting [16, 41] or synergy in training supervised segmentation models [44, 17]. However, due to the scarcity of public datasets, existing adaptions mostly focus on structural medical images like CT and MRI, ignoring molecular PET images.
Annotation-Efficient Learning. Despite the outstanding performance in many medical image segmentation tasks [9, 36], the success of existing deep neural networks heavily relies on the massive radiologist-examined labeled data. Obtaining pixel-wise annotation is time-consuming and expertise-demanding, which motivates annotation-efficient learning strategies to train deep models with limited annotations [2, 13, 32]. Despite recent efforts of large-scale medical datasets [39, 1], these datasets are the integration of existing public datasets with different annotation quality. While some labels may be noisy with mistakes and could potentially hinder the learning procedure, it is important to develop noise-robust learning strategies.
@@ 3 Methodology
6:   Input prompt $p_{i}$ $\Leftarrow$ PromptGenerate($Y$, $\hat{Y}_{i-1}$)
@@ 3.1 Network Architecture
Motivated by the recent advance of segmentation foundation models for natural images [15], we reformulate a holistic 3D structure in SegAnyPET to capture the spatial information directly from volumetric images for universal segmentation. Specifically, SegAnyPET consists of three main components, including an image encoder $\mathcal{F}_{IE}$, a prompt encoder $\mathcal{F}_{PE}$, and a mask decoder $\mathcal{F}_{D}$. The image encoder aims to transform input images into discrete embeddings, and the prompt encoder converts input prompts into compact embeddings by combining fixed positional encoding and adaptable prompt-specific embeddings. After that, the mask decoder receives the extracted information from both the image encoder and the prompt encoder and incorporates prompt self-attention and cross-attention in two directions for prompt-to-image and image-to-prompt attention to update the image embeddings and prompt embeddings. The processed feature map is then up-sampled and passes through a multi-layer perception to generate the output segmentation masks. Details of the network components are shown in the Appendix. Given an input image $X$ and spatial prompt $p$, the output segmentation mask $\hat{Y}$ can be formally expressed as:
[…]
where $\mathbf{O}_{H\times W\times D}$ is an all-zero matrix as the pre-defined initial segmentation mask. In the training loop, new point prompts are generated from error regions based on the previous segmentation mask to simulate the manual interactive segmentation workflow.
@@
SegResNet [25] Auto 92.28 85.56 82.94 88.21 86.55 87.11
SwinUNETR [6] Auto 93.03 89.23 83.62 89.03 87.93 88.57
nnUNet [9] Auto 93.16 90.12 86.60 90.96 88.32 89.83
STUNet [8] Auto 93.15 90.11 85.60 90.38 88.83 89.61
Table 1: Comparison of DSC performance [%] with state-of-the-art segmentation foundation models for zero-shot interactive segmentation and task-specific models for training-based automatic segmentation from PET images. N denotes the count of slices containing the target object (N ranges from 20 to 50 in our task).
@@ 3.2 Cross Prompting Confident Learning
Due to the dependency of domain knowledge for annotating medical images, acquiring large amounts of high-quality labeled data is infeasible. A prevalent approach to mitigating this issue is collecting additional labeled data with varying annotation qualities like crowdsourcing from non-experts or using model-generated labels. Given that some annotations may be fraught with noise or inaccuracies, which could impede the learning process, it is imperative to devise noise-robust learning strategies to efficiently learn from both expert-examined high-quality annotations (termed as HQ Set) and low-quality annotations with possible mistakes (termed as LQ Set). To issue the challenge of discrepant annotation quality, we adopt a cross prompting confident learning (CPCL) strategy with a cross prompting consistency similarity regularization and an uncertainty-guided self-rectification process. Figure 3 depicts the workflow of CPCL. The utilization of LQ set can be mainly divided into the following two parts.
@@ 3.3 Overall Training Procedure
Aorta Lung-$L$ Lung-LR Lung-UL Lung-UR Lung-MR Prostate Avg
@@
Table 2: Generalization performance to training invisible organs with comparison to state-of-the-art segmentation foundation models for zero-shot interactive segmentation from PET images.
4 Experiments
4.1 Datasets and Experimental Setup
@@
We conduct comprehensive experiments on two different PET datasets. The first dataset is the proposed PETS-5k dataset, which consists of 5,731 three-dimensional whole-body PET images with the annotation of 5 target organs. Among the dataset, a subset of 100 images are with high-quality annotations verified by domain experts, which is split into 40 cases serving as HQ training set and 60 cases serving as internal test set. The remaining 5,631 images are noisy-annotated with possible mislabeled or unlabeled pixels serving as the LQ training set. To evaluate the generalization performance for universal segmentation, we annotate 7 additional training invisible organs from the internal test set. The second dataset is AutoPET-Organ dataset extended from the AutoPET dataset [4] consists of 1,014 three-dimensional whole-body PET images with corresponding annotation of tumor lesions. To evaluate the generalization ability, we further annotate a small subset with 100 cases with all 12 organs as an external test set. Following the setting for promptable segmentation, the input prompts are simulated based on the ground truth mask with random perturbations. We compare proposed SegAnyPET with state-of-the-art segmentation foundation models [15, 21, 34] and training-based task-specific models [25, 9, 6, 8]. More information of the datasets and implementation details are shown in the Appendix.
4.2 Experiments